OpenStudy (anonymous):

dont get how to do this! help Write the equation y - 4 = -2(x - 5) in slope-intercept form. A. y = -2x + 14 B. y = -2x + 5 C. y = -2x + 9 D. y = -2x + 6

OpenStudy (anonymous):

A y-4=-2(x-5) y=-2x+14

OpenStudy (anonymous):

oh ok that makes more sense. thank you